# FedADMM-InSa: An Inexact and Self-Adaptive ADMM for Federated Learning

### v1

@@ 3.1 Server’s Aggregation with Memory
From (2.12), it is evident that in the vanilla FedADMM, the update of the server’s global model parameters $z^{k+1}$ depends solely on the local parameters updated by clients in the $(k+1)$-th communication round. Due to the uncertainty arising from heterogeneous local updates, the global model parameters may undergo significant fluctuations. To address this challenge, we draw inspiration from exponential moving average methods used in several fields (see e.g., [2, 5, 6]) and propose an improved method for updating global model parameters $z^{k+1}$ as follows:
@@
where $\delta>0$ controls the trade-off between stability and responsiveness in the server update process. The proposed server update strategy (3.1) improves the stability of aggregation by balancing the current and previous global model parameters. This ensures the robustness of the global model in the face of variable clients’ local updates. On the one hand, it allows for the memorization of $z^{k}$, thereby minimizing the sensitivity to outliers and leading to a more stable global model. On the other hand, the strategy maintains adaptability to the changing $\hat{z}^{k+1}$, ensuring that the global model responds appropriately to evolving local updates. Meanwhile, the proposed server update strategy (3.1) can be interpreted by augmenting the vanilla $z$-subproblem (2.8c) with an additional proximal term, given by
@@ 3.2 An Inexact Version of FedADMM
where $\tilde{\beta}_{i}={\beta_{i}}/{c_{i}}$. It is evident that a smaller $\tilde{\beta}_{i}$ leads to a higher permissible value of $\sigma_{i}$, indicating that the subproblem can be solved with less precision. In practical scenarios, each client can (locally) convexify its local loss function by incorporating a regularization term. Our inexactness criterion can also be applied to the non-convex problems by using a $\tilde{\beta_{i}}>0$ in (3.9). Subsequent numerical experiments in Section 4, covering both strongly convex and non-convex problems, also underscore the feasibility of our inexactness criterion (3.7). It is noteworthy that our inexactness criterion can be assessed by each client autonomously based on its present model and penalty parameter, and it can be seamlessly executed during iterations. There is no requirement to predefine any empirically perceived constant accuracy. This gives each client the flexibility to solve its subproblem inexactly, aligning with its distinct non-IID data and computational resources. Consequently, this approach effectively mitigates the potential risk of the straggler effect that may arise with resource-constrained clients. Overall, these attributes make the inexactness criterion (3.7) straightforward to implement and more likely to result in local computational savings.
FedADMM-In: An inexact FedADMM algorithm.
@@
8:           Update $\hat{u}_{i}^{k}\coloneqq\hat{u}_{i}^{k}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}(\hat{u}_{i}^{k},\lambda_{i}^{k},z^{k})$.
@@
16:        Send $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and ${\beta_{i}}$ to the server.
17:     end for
18:     Server side:
19:     Obtain $\hat{z}^{k+1}=\frac{1}{\sum_{i=1}^{m}\alpha_{i}\beta_{i}}\sum_{i=1}^{m}\alpha_{i}(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$.
20:     Update $z^{k+1}=\frac{1}{1+\delta}\hat{z}^{k+1}+\frac{\delta}{1+\delta}z^{k}$.
21:  end for
Convergence analysis.
@@
The detailed proof is presented in Appendix A.1, which is inspired by [11, Sec. 3.3]. Although the convergence proof is conducted under the strong convexity assumption, our algorithm with the proposed inexactness criterion can also be applied to the non-convex cases by using a $\tilde{\beta_{i}}>0$ in (3.9). Moreover, the effectiveness of the algorithm with partial client participation in both strongly convex and non-convex cases is validated by the numerical experiments in Section 4.
@@ 3.3 Self-Adaptive Penalty Parameter $\beta_{i}^{k}$
Self-adaptive scheme.
@@
where $\mu,\tau>1$ are parameters to choose, and a typical choice can be $\mu=20$ and $\tau=2$. The essence of the self-adaptive penalty parameter scheme (3.15) lies in dynamically balancing the primal residual $p_{i}^{k}$ and the dual residual $d_{i}^{k}$. For instance, $d_{i}^{k}>\mu p_{i}^{k}$ indicates an excessive dual residual $d_{i}^{k}$. Following the definition of $d_{i}^{k}$ in (3.14b), this suggests that client $i$’s updated local model parameters $u_{i}^{k+1}$ deviate significantly from the current global model parameters $z^{k}$. Consequently, client $i$ is prompted to increase its penalty parameter $\beta_{i}^{k}$ to impose a more substantial penalty on the constraint. Conversely, a reduction in $\beta_{i}^{k}$ is recommended for client $i$ to mitigate the primal residual. Finally, the iterative steps of FedADMM with an adaptive $\beta_{i}^{k}$ are given by
@@
Compatibility with the inexactness criterion.
@@
FedADMM-InSa: An inexact and self-adaptive FedADMM.
@@
8:           Update $\hat{u}_{i}^{k}\coloneqq\hat{u}_{i}^{k}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}(\hat{u}_{i}^{k},\lambda_{i}^{k},z^{k})$.
@@
25:        Send $(\beta_{i}^{k}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and ${\beta_{i}^{k}}$ to the server.
26:     end for
27:     Server side:
28:     Obtain $\hat{z}^{k+1}=\frac{1}{\sum_{i=1}^{m}\alpha_{i}\beta_{i}^{k}}\sum_{i=1}^{m}\alpha_{i}(\beta_{i}^{k}u_{i}^{k+1}-\lambda_{i}^{k+1})$.
29:     Update $z^{k+1}=\frac{1}{1+\delta}\hat{z}^{k+1}+\frac{\delta}{1+\delta}z^{k}$.
30:  end for
@@ 4 Experimental Results
In this section, we conduct extensive numerical tests to demonstrate the improved performance of our proposed FedADMM-In (Algorithm 2) and FedADMM-InSa (Algorithm 3) in comparison to the vanilla FedADMM (Algorithm 1). We first introduce the experimental settings and the implementation details used in our experiments. Then, we present the simulation results and analysis.
@@ 4.1 Setups
Example 1: Linear regression with a synthetic dataset.
@@
Example 2: Image classification with the MNIST dataset.
In the second example, we address the image classification problem using fully-connected neural networks with the MNIST dataset [20]. The MNIST dataset contains images of handwritten digits. The samples in this dataset are $28\times 28$ grayscale images with handwritten digits from 0 to 9 in the center. We select a training set of 10,000 examples and a test set of 1,000 examples. The number of clients is set to $m=100$. For the neural networks, each client uses a fully-connected neural network with two hidden layers, each comprising 200 neurons, incorporating the cross-entropy loss and the ReLU activation function.
Non-IID data separation.
To test the performance of the algorithms on non-IID data, we separate datasets in the following way. In both examples, there are in total $N=10,000$ training samples and $m=100$ clients. In Example 1, we first shuffle the training samples and then distribute them evenly among all clients. In Example 2, we first sorted the data by labels and then divided them into several shards, each shard containing images with the same label. Then, for 100 clients, the training dataset with 10,000 samples is divided into 200 shards of size 50, and each client receives 2 shards. This means each client has at most two kinds of labels in their dataset, corresponding to a pathological non-IID scenario.
FL training parameters.
In each communication round, the server uniformly samples 10% of the clients to perform the local training. During the local training, the selected clients use full-batch gradient descent with a learning rate of 0.01. For the vanilla FedADMM algorithm, unless otherwise indicated, the clients perform fixed $E=10$ steps of the full-batch gradient descent. For FedADMM-In and FedADMM-InSa algorithms, the number of the full-batch gradient descent steps is controlled by the inexactness criterion, and we constrain the maximum number of gradient descent steps in each communication round to 10. For the inexactness criterion, we set $\tilde{\beta}_{i}$ to the same value as ${\beta_{i}}$ and replace $u_{i}^{k}$ with $z^{k}$ for all the tests. We use the weight coefficient $\alpha_{i}=0.01$ because the amount of data per client is the same in both examples. We choose $\delta=0.01$ for the server aggregation. Finally, important parameters used in the tests are summarized in Table 1.
Table 1: Dataset parameters and FL training parameters.
Parameters
@@
Training set size 10000 10000
Test set size - 1000
Data dimension 1000 $28\times 28$
Communication rounds $K$ 100 200
Number of clients $m$ 100 100
Client subset size $|\mathcal{M}^{k}|$ 10 10
Client learning rate $\eta_{i}$ 0.01 0.01
4.2 Evaluation of FedADMM-In (Algorithm 2) with Example 1
In this subsection, we present the comparison results of our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1) with Example 1. The results are plotted in Figure 1. In this experiment, we test scenarios with $\beta_{i}=1$ and $\beta_{i}=2$. In the case of $\beta_{i}=2$, as shown in Figure 0(a), FedADMM-In achieves lower loss values from the beginning and maintains this advantage throughout the experiment. For $\beta_{i}=1$, although FedADMM-In initially exhibits a slower loss decline than FedADMM, it reaches a similar loss value to FedADMM after about 30 rounds. Meanwhile, as illustrated in Figure 0(b) and Figure 0(c), FedADMM-In requires significantly fewer local gradient descent steps than FedADMM. These results indicate that our proposed FedADMM-In algorithm can achieve comparable or even superior training outcomes while reducing the computational load on clients.
@@ 4.2 Evaluation of FedADMM-In (Algorithm 2) with Example 1
(b) Client 1’s local update steps.
(c) Client 1’s local update steps.
Figure 1: Comparison results of our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1) with Example 1. For the vanilla FedADMM, the clients perform fixed $E=10$ local full-batch gradient descent steps. For our FedADMM-In, the number of full-batch gradient descent steps is controlled by the inexactness criterion (3.7).
4.3 Evaluation of FedADMM-In (Algorithm 2) with Example 2
In this subsection, we conduct the comparison tests between our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1) with Example 2 and present the results in Figure 2. In both sets of experiments conducted with penalty parameters $\beta_{i}=1$ and $\beta_{i}=2$, our proposed FedADMM-In consistently demonstrates superior performance. This is evidenced by achieving higher prediction accuracy and lower loss after the same 200 communication rounds, as illustrated in Figure 1(a) and Figure 1(b), respectively. Additionally, we visualize the number of local gradient descent steps performed by a specific client (client 1) in each communication round, represented in Figure 1(c) and Figure 1(d). It is evident that employing our proposed FedADMM-In algorithm allows clients to execute fewer gradient descent steps as well as ensure training efficiency. This significantly reduces the potential waste of valuable computational resources in FL. However, it is noteworthy from Figure 2 that when $\beta_{i}=2$, the performance of both algorithms deteriorates compared to the case of $\beta_{i}=1$. This discrepancy reflects the substantial impact of the initial pre-selected penalty parameters on the training performance. To address this issue, we propose the FedADMM-InSa algorithm, the performance of which is validated in the next subsection.
(a) Test accuracy.
(b) Training loss.
(c) Client 1’s local update steps ($\beta_{i}=1$).
(d) Client 1’s local update steps ($\beta_{i}=2$).
Figure 2: Comparison results of our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1) with Example 2. For the vanilla FedADMM, the clients perform fixed $E=10$ local full-batch gradient descent steps. For our FedADMM-In, the number of the full-batch gradient descent steps is controlled by the inexactness criterion (3.7).
@@ 4.3 Evaluation of FedADMM-In (Algorithm 2) with Example 2
4.4 Evaluation of FedADMM-InSa (Algorithm 3) with Example 2
In this subsection, we present the comparison results of our FedADMM-InSa (Algorithm 3) with the other two algorithms, namely our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1). For FedADMM-InSa, we set the initial penalty parameter $\beta_{i}^{0}=2$ and update it by the scheme (3.15) with $\mu=20$ and $\tau=2$. For both FedADMM-In and FedADMM, the penalty parameter $\beta_{i}=2$ is fixed throughout the training. For the vanilla FedADMM, we test it with three different local full-batch gradient descent steps, i.e., $E=2,5,10$. The results are plotted in Figure 3 and summarized in Table 2. From Figure 3, it is evident that our FedADMM-InSa algorithm, incorporating a self-adaptive penalty parameter (3.15), significantly outperforms the other two algorithms. As depicted in Figure 2(a), under the same number of communication rounds, FedADMM-InSa exhibits significantly higher prediction accuracy compared to the other two algorithms. In other words, FedADMM-InSa can achieve a given target accuracy with fewer communication rounds than the other two algorithms. This is also validated by the fastest decreasing training loss of FedADMM-InSa shown in Figure 2(b). Figure 2(d) illustrates that for a specific client (client 1), when using FedADMM-InSa, it performs less number of gradient descent steps compared with FedADMM-In and FedADMM ($E=10$), especially in the later communication rounds. This reduction in gradient descent steps is further quantified in Table 2, where compared to the baseline FedADMM ($E=10$) algorithm, our proposed FedADMM-In and FedADMM-InSa algorithms reduce the number of gradient descent steps by 49.8% and 64.3%, respectively, leading to a substantial improvement in training efficiency. Although FedADMM ($E=2$) has the lowest total number of local gradient descent steps, the value of $E$ has to be chosen empirically in advance. Without knowing the empirical information of the FL system, it is difficult to find an appropriate $E$ in advance. Meanwhile, the decline in test accuracy from FedADMM ($E=2$) to FedADMM ($E=10$) is likely attributed to the pathological non-IID datasets, further emphasizing the challenge of empirically choosing an appropriate $E$ in advance. Notably, our proposed FedADMM-InSa algorithm outperforms FedADMM ($E=2$) in terms of test accuracy, achieving a balance between improving training performance and reducing computational load. Finally, in Figure 2(c), the variation of the average penalty parameter across all clients is plotted. It can be observed that FedADMM-InSa dynamically adjusts the average penalty parameter from an initial value of two to around one. This demonstrates that FedADMM-InSa can adaptively correct the initially improperly chosen penalty parameters, highlighting the robustness of our proposed algorithm.
(a) Test accuracy.
(b) Training loss.
(c) Average of all clients’ penalty parameters.
(d) Client 1’s local update steps.
Figure 3: Comparison results of our FedADMM-InSa (Algorithm 3), our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1) with Example 2. For our FedADMM-InSa, the initial penalty parameter $\beta_{i}^{0}=2$ and updates by scheme (3.15) with $\mu=20$ and $\tau=2$. For the vanilla FedADMM, the client’s local full-batch gradient descent steps $E=2,5,10$. For both FedADMM-In and FedADMM, the penalty parameter is fixed as $\beta_{i}=2$ throughout the training.
Table 2: Detailed comparison results of the tests presented in Figure 3. The results in the steps reduction column use the vanilla FedADMM ($E=10$) as the baseline.
Algorithm Test accuracy Training loss Local update steps Steps reduction
FedADMM ($E=2$) 81.6% 0.528 4000 -
FedADMM ($E=5$) 71.9% 0.762 10000 -
FedADMM ($E=10$) 62.6% 0.955 20000 0.0%
FedADMM-In (ours) 70.9% 0.730 10036 49.8%
FedADMM-InSa (ours) 87.8% 0.325 7139 64.3%

### v3

@@ 3.1 Server’s Aggregation with Memory
From (2.12), it is evident that in the vanilla FedADMM, the update of the server’s global model parameters $z^{k+1}$ depends solely on the local parameters updated by clients in the $(k+1)$-th communication round. Due to the uncertainty arising from heterogeneous local updates, the global model parameters may undergo significant fluctuations. To address this challenge, we draw inspiration from exponential moving average methods used in several fields (see e.g., (Awheda and Schwartz, 2016; Cai et al., 2021; Dinh et al., 2020)) and propose an improved method for updating global model parameters $z^{k+1}$ as follows:
@@
where $\delta>0$ controls the trade-off between stability and responsiveness in the server update.
The proposed server update strategy (3.1) improves the stability of aggregation by balancing the current and previous global model parameters. This ensures the robustness of the global model in the face of variable clients’ local updates. On the one hand, it allows for the memorization of $z^{k}$, thereby minimizing the sensitivity to outliers and leading to a more stable global model. On the other hand, the strategy maintains adaptability to the changing $\hat{z}^{k+1}$, ensuring that the global model responds appropriately to evolving local updates.
Meanwhile, the proposed server update strategy (3.1) can be interpreted by augmenting the vanilla $z$-subproblem (2.8c) with an additional proximal term, given by
@@ 3.2 An Inexact Version of FedADMM
where $\tilde{\beta}_{i}={\beta_{i}}/{c_{i}}$. It is evident that a smaller $\tilde{\beta}_{i}$ leads to a higher permissible value of $\sigma_{i}$, indicating that the subproblem can be solved with less precision. In practical scenarios, each client can (locally) convexify its local loss function by incorporating a regularization term. Our inexactness criterion can also be applied to the non-convex problems by using a $\tilde{\beta_{i}}>0$ in (3.9). Subsequent numerical experiments in Section 4, covering both strongly convex and non-convex problems, also underscore the feasibility of our inexactness criterion (3.7).
It is noteworthy that our inexactness criterion can be assessed by each client autonomously based on its present model and penalty parameter, and it can be seamlessly executed during iterations. There is no requirement to predefine any empirically perceived constant accuracy. This gives each client the flexibility to solve its subproblem inexactly, aligning with its distinct non-IID data and computational resources. Consequently, this approach effectively mitigates the potential risk of the straggler effect that may arise with resource-constrained clients. Overall, these attributes make the inexactness criterion (3.7) straightforward to implement and more likely to result in local computational savings.
FedADMM-In: An inexact FedADMM algorithm
@@
8:           Update $\hat{u}_{i}^{k}\coloneqq\hat{u}_{i}^{k}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}(\hat{u}_{i}^{k},\lambda_{i}^{k},z^{k})$ for one epoch.
@@
16:     end for
17:     Server side:
18:     Obtain $\hat{z}^{k+1}=\frac{1}{\sum_{i=1}^{m}\alpha_{i}\beta_{i}}\sum_{i=1}^{m}\alpha_{i}(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$.
19:     Update $z^{k+1}=\frac{1}{1+\delta}\hat{z}^{k+1}+\frac{\delta}{1+\delta}z^{k}$.
20:  end for
Convergence analysis
@@
The detailed proof is presented in Appendix A.1, which is inspired by (Glowinski et al., 2022, Sec. 3.3). Although the convergence proof is conducted under the strong convexity assumption, our algorithm with the proposed inexactness criterion can also be applied to the non-convex cases by using a $\tilde{\beta_{i}}>0$ in (3.9). Moreover, the effectiveness of the algorithm with partial client participation in both strongly convex and non-convex cases is validated by the numerical experiments in Section 4.
@@ 3.3 Self-Adaptive Penalty Parameter $\beta_{i}^{k}$
Self-adaptive scheme
@@
where $\mu,\tau>1$ are parameters to choose.
The essence of the self-adaptive penalty parameter scheme (3.15) lies in dynamically balancing the primal residual $p_{i}^{k}$ and the dual residual $d_{i}^{k}$. For instance, according to the definition of $d_{i}^{k}$ in (3.14b), $d_{i}^{k}>\mu p_{i}^{k}$ suggests that client $i$’s updated local model parameters $u_{i}^{k+1}$ deviate significantly from the current global model parameters $z^{k}$. Consequently, client $i$ is prompted to increase its penalty parameter $\beta_{i}^{k}$ to impose a more substantial penalty on the constraint. Conversely, a reduction in $\beta_{i}^{k}$ is recommended for client $i$ to mitigate the primal residual.
Finally, the iterative steps of FedADMM with an adaptive $\beta_{i}^{k}$ are given by
@@
Compatibility with the inexactness criterion
@@
FedADMM-InSa: An inexact and self-adaptive FedADMM
@@
8:           Update $\hat{u}_{i}^{k}\coloneqq\hat{u}_{i}^{k}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}(\hat{u}_{i}^{k},\lambda_{i}^{k},z^{k})$ for one epoch.
@@
25:     end for
26:     Server side:
27:     Obtain $\hat{z}^{k+1}=\frac{1}{\sum_{i=1}^{m}\alpha_{i}\beta_{i}^{k}}\sum_{i=1}^{m}\alpha_{i}(\beta_{i}^{k}u_{i}^{k+1}-\lambda_{i}^{k+1})$.
28:     Update $z^{k+1}=\frac{1}{1+\delta}\hat{z}^{k+1}+\frac{\delta}{1+\delta}z^{k}$.
29:  end for
@@ 4 Experimental Results
In this section, we conduct extensive numerical tests to demonstrate the improved performance of our proposed FedADMM-In (Algorithm 2) and FedADMM-InSa (Algorithm 3) in comparison to the vanilla FedADMM ((Zhou and Li, 2023)) and FedAvg (McMahan et al., 2017). We first introduce the experimental settings and the implementation details used in our experiments. Then, we present the simulation results and analysis.
@@ 4.1 Setups
We conducted experiments on three combinations of datasets and models, covering both cross-device (many clients with few data points per client) and cross-silo (few clients with many data points per client) scenarios (Kairouz et al., 2021). The details of the experiments are elaborated below and also summarized in Table 1.
Example 1: Linear regression with a synthetic dataset
@@
Example 2: Image classification with the MNIST dataset
In the second example, we address the image classification problem using convolutional neural networks (CNN) with the MNIST dataset (LeCun et al., 1998). The MNIST dataset contains images of handwritten digits. It has 60,000 training data and 10,000 test data. The samples in this dataset are $28\times 28$ grayscale images with handwritten digits from 0 to 9 in the center. In this example, each client uses a CNN with two convolutional layers, comprising 32 and 64 channels respectively, as the same to that in (McMahan et al., 2017).
Example 3: Image classification with the CIFAR-10 dataset
In the third example, we test the image classification of the CIFAR-10 dataset (Krizhevsky and Hinton, 2009). This dataset contains color images of size $32\times 32$ with 10 categories. There are 50,000 training data and 10,000 test data. For the neural networks, each client uses a ResNet-20 (He et al., 2016), which consists of 20 stacked weighted layers.
Non-IID data separation
To test the performance of the algorithms on non-IID data, we separate datasets in the following way. In Example 1, we first shuffle the training samples and then distribute them evenly among all clients. In Example 2 and 3, we first sorted the data by labels and then divided them into several shards, each shard containing images with the same label. Then, we assign the data to make sure that each client has at most two kinds of labels in Example 2 and five kinds of labels in Example 3, corresponding to pathological non-IID scenarios.
FL training parameters
In each round, the server uniformly samples 20% of the clients to perform the local training. For the vanilla FedADMM and FedAvg algorithms, the clients perform fixed epochs of training using stochastic gradient descent. For the FedADMM-In and FedADMM-InSa algorithms, the number of epochs is controlled by the inexactness criterion, which is evaluated after each epoch. We constrain the maximum number of epochs per round to be the same as the fixed epochs in the vanilla FedADMM. For the inexactness criterion, we set $c_{i}=0.01$ in Example 1 and 2, and $c_{i}=0.001$ in Example 3. In all examples, we replace $u_{i}^{k}$ with $z^{k}$ and set $\delta=0.01$ for the server aggregation. For FedADMM-InSa, the adaptive penalty scheme uses $\mu=5$ and $\tau=2$ in all examples. Finally, important parameters used in the tests are summarized in Table 1.
Table 1: Datasets, models, and parameters of three examples.
@@
Example 3
Image classification
Dataset Synthetic MNIST CIFAR-10
Model Linear CNN ResNet
Training set size 50000 60000 50000
Test set size - 10000 10000
Data dimension 5000 $28\times 28\times 1$ $32\times 32\times 3$
Number of clients 200 200 10
Data per client 250 300 5000
Active clients per round 20% 20% 20%
Client learning rate 0.001 0.01 0.001
Batch size 50 50 500
Epochs per round 20 20 10
Communication rounds 300 300 300
Table 2: Comparison results of different examples. The columns showing epoch reduction use the vanilla FedADMM and FedAvg as the baseline.
Example 1
Linear regression
Example 2
MNIST, CNN
Example 3
CIFAR-10, ResNet-20
$\beta_{i}$ Algorithm Loss
Epoch
reduction
Accuracy
Epoch
reduction
Accuracy
Epoch
reduction
- Fedavg 1.64 - 96.0% - 36.3% -
0.1 FedADMM 1.64 - 98.6% - 51.1% -
FedADMM-In 1.63 94.3% 98.5% 92.4% 41.0% 71.5%
FedADMM-InSa 1.51 20.3% 97.8% 66.9% 53.1% 14.4%
1 FedADMM 1.53 - 97.7% - 52.9% -
FedADMM-In 1.55 58.5% 97.6% 58.7% 51.6% 9.4%
FedADMM-InSa 1.51 18.8% 97.9% 61.0% 52.1% 9.4%
2 FedADMM 1.51 - 96.2% - 46.7% -
FedADMM-In 1.51 19.3% 96.7% 41.7% 45.7% 2.1%
FedADMM-InSa 1.51 16.2% 97.7% 57.9% 50.4% 6.9%
5 FedADMM 1.51 - 92.5% - 39.5% -
FedADMM-In 1.51 3.9% 93.4% 31.1% 39.3% 0.0%
FedADMM-InSa 1.51 12.5% 97.9% 55.9% 51.5% 4.0%
10 FedADMM 1.51 - 89.7% - 34.0% -
FedADMM-In 1.51 0.9% 90.2% 21.1% 37.3% 0.0%
FedADMM-InSa 1.51 6.8% 97.5% 55.7% 49.8% 4.8%
4.2 Results Analysis
In this subsection, we present the comparison results of our FedADMM-In and FedADMM-InSa algorithms with benchmark FedADMM and FedAvg algorithms. For ADMM-based algorithms, we show results with different values of $\beta_{i}\in\{0.1,1,2,5,10\}$. The detailed results are summarized in Table 2 and elaborated below.
4.2.1 Evaluation of FedADMM-In
In this subsection, we compare our FedADMM-In algorithm with the vanilla FedADMM and FedAvg algorithms. The results with the penalty parameter $\beta_{i}=1$ are plotted in Figure 1, while other results are summarized in Table 2.
As shown in Figures 1(a)-1(c), FedADMM-In achieves superior results compared to FedAvg and comparable results to FedADMM across all three examples. Furthermore, Figures 1(d)-1(f) demonstrate that FedADMM-In allows clients to execute fewer local epochs compared to the other two algorithms while maintaining training performance. As detailed in Table 2, FedADMM-In achieves an average local epoch reduction of 35.4%, 49.0%, and 16.6% in Examples 1-3, respectively, leading to a substantial improvement in computational savings. These results indicate that our proposed FedADMM-In algorithm can achieve comparable or superior training outcomes while reducing the computational load on clients, significantly mitigating the potential waste of valuable computational resources in FL.
(a) Linear regression.
(b) MNIST with CNN.
(c) CIFAR-10 with ResNet.
(d) Linear regression.
(e) MNIST with CNN.
(f) CIFAR-10 with ResNet.
Figure 1: Comparison results of FedADMM-In in three examples.
4.2.2 Evaluation of FedADMM-InSa
In this subsection, we present comparison results of our FedADMM-InSa algorithm with the vanilla FedADMM and FedAvg algorithms. The results for the three examples are presented in Figures 2-4 and are also summarized in Table 2.
Results of Example 1 (Linear regression)
Figure 2 shows the comparison results of our FedADMM-InSa with the vanilla FedADMM and FedAvg algorithms. We plot scenarios with $\beta_{i}\in\{0.1,1\}$, and other results can be seen in Table 2. It can be seen from Figure 2(a) that our FedADMM-InSa consistently achieves the lowest training loss. For FedADMM, it performs better than FedAvg when $\beta_{i}=1$. However, for $\beta_{i}=0.1$, FedADMM only decreases to a loss value similar to that of FedAvg, while our FedADMM-InSa reaches the lowest loss value. The average $\beta_{i}$ of clients are plotted in Figure 2(b). For FedADMM-InSa, despite different initial $\beta_{i}$ values, they converge to a similar value around three. Additionally, Figure 2(c) presents the number of average local epochs of active clients in each round. It is evident that FedADMM-InSa requires fewer local epochs than FedADMM and FedAvg. In the later training rounds, the local epochs saturate probably because the algorithm has approached the local minimum.
@@ 4.2.2 Evaluation of FedADMM-InSa
(b) Average penalty parameters.
(c) Average local epochs.
Figure 2: Comparison results of FedADMM-InSa in Example 1 (Linear regression).
Results of Example 2 (MNIST with CNN)
Figure 3 shows the comparison results on the MNIST dataset using a CNN model. Similar to Example 1, the performance of FedADMM varies significantly with different $\beta_{i}$ values. In the two cases plotted with $\beta_{i}\in\{5,10\}$, FedADMM performs even worse compared to FedAvg, highlighting its sensitivity to the choice of $\beta_{i}$. Conversely, our proposed FedADMM-InSa consistently achieves superior performance, demonstrating lower training loss and higher test accuracy across all algorithms.
Figure 3(c) illustrates the change in average penalty parameters of all clients. It can be seen that FedADMM-InSa maintains an effective penalty parameter adaptation process, contributing to its robust performance. Meanwhile, Figure 3(d) shows the average local epochs. FedADMM-InSa requires much fewer local epochs on average compared to FedADMM, indicating a more efficient training process. This efficiency is critical in FL scenarios, as the client can complete local training faster without compromising the global model accuracy.
(a) Training loss.
(b) Test accuracy.
(c) Average penalty parameters.
(d) Average local epochs.
@@
Figure 3: Comparison results of FedADMM-InSa in Example 2 (MNIST with CNN).
Results of Example 3 (CIFAR-10 with ResNet-20)
In Figure 4, the performance of different algorithms on the CIFAR-10 dataset using a ResNet-20 model is presented. Similar to the previous two examples, FedADMM-InSa achieves the best performance, as evidenced by the lowest training loss in Figure 4(a) and the highest test accuracy in Figure 4(b). The average local epochs required by FedADMM-InSa, as shown in Figure 4(d), are still slightly lower than those conducted by FedADMM and FedAvg. The penalty parameter updates in Figure 4(c) illustrate the adaptive capability of FedADMM-InSa, contributing to its overall effectiveness. In summary, the comparison results across all three examples consistently demonstrate that our FedADMM-InSa outperforms both FedAvg and FedADMM in terms of test accuracy, workload reduction, and robustness across various penalty parameters.
(a) Training loss.
(b) Test accuracy.
(c) Average penalty parameters.
(d) Average local epochs.
Figure 4: Comparison results of FedADMM-InSa in Example 3 (CIFAR-10 with ResNet).
4.2.3 Evaluation of the Self-Adaptive Penalty Parameter Scheme (3.15)
As observed in the three examples above, the pre-selected penalty parameter $\beta_{i}$ plays a crucial role in the performance of the FedADMM algorithm. For example, it performs better with a larger $\beta_{i}$ in Example 1, but with a smaller $\beta_{i}$ in Examples 2 and 3. In this subsection, we compare the results of FedADMM and FedADMM-InSa under different values of $\beta_{i}\in\{0.1,1,2,5,10\}$ using Example 2. The results of the other two examples are similar and omitted to save space.
It is clear from Figure 5 that our proposed FedADMM-InSa (bottom row) demonstrates robustness across different values of $\beta_{i}$, while FedADMM (top row) only performs well under certain values. As shown in Figure 5(f), our proposed adaptive scheme dynamically adjusts the penalty parameter and ensures it converges to stable values, allowing our algorithm to maintain robust performance across different initial penalty parameter values. This adaptive mechanism prevents over-penalization, which can hinder convergence, and under-penalization, which can lead to insufficient coordination among clients. As a result, as shown in Figure 5(d) and Figure 5(e), FedADMM-InSa achieves lower training loss and higher test accuracy. Moreover, FedADMM-InSa also requires fewer local epochs compared to FedADMM, as detailed in Table 2. Overall, our adaptive penalty parameter scheme (3.15) enhances the flexibility and robustness of our FedADMM-InSa algorithm, making it a superior choice for FL applications where training efficiency and computational resources are critical considerations.
(a) Training loss.
(b) Test accuracy.
(c) Average penalty parameters.
(d) Training loss.
(e) Test accuracy.
(f) Average penalty parameters.
Figure 5: Comparison of FedADMM and FedADMM-InSa with different values of $\beta_{i}$ in Example 2. FedADMM (top row) uses fixed penalty parameters. Our FedADMM-InSa (bottom row) utilizes adaptive penalty parameter scheme (3.15).